OpenStudy (anonymous):

multiplying rational expressions 3x/x+2 (times) X/x-1

OpenStudy (anonymous):

3x^2/x^2+x-2

OpenStudy (anonymous):

Hey there beautiful ;) What's your name?

OpenStudy (anonymous):

aweee arent you sweet.. :) umm if we could please keep it about math on here.. thanks

OpenStudy (anonymous):

\[\frac{3x}{x+2}\times \frac{x}{x-1}=\frac{3x\times x}{(x+2)(x-1)}\] \[=\frac{3x^3}{x^2+x-1}\]

OpenStudy (anonymous):

thanks sooo muchhh satellite73